In 2005, for each of the 14 million people present in a country, 0.028 were born and 0.008 died during the year. Using exponential equation, the number of people present in 2015 is predicted as:
a. 25 millions
b. 17 millions
c. 20 millions
d. 18 millions

Grade:12

1 Answers

Beenu Mathew
7 years ago
HI,
u would have seen this weird formula in organisms n populations chapter n ie
Nt= No * ert
where
r is the
intrinsic rate of natural increase and
t- time period
r can be calculated as follows r=b-d
ie 
birthrate – death rate
in this q 
b=0.028
d= 0.008 , therefore b-d=.020
Nt = 14 million * e.020 * 10(by the way the time period is 10 years ie from 2005-2010)
plug in the values and after solving you will get
Nt=17 million
 
HOPE THIS HELPS
JEHOYESJIRA